the angular frequency 'w' of a 200v A.C source is 3.14rads–¹ .if the peak currwnt is 5A calculate the power decipated
1
Expert's answer
2020-05-21T13:46:18-0400

"P=\frac{U_{max}\times I_{max}}{2}=\frac{200\times 5}{2}=500W"

Answer: P = 500W